31. e^ln(3) =
s. 1
b. 2
c. 0
d. e
Can someone help me out with this?
Eddie


Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

Determine which two equations represent parallel lines.
(a) y = 5/3x + 3
(b) y = 3/5x + 6
(c) y = 3/5x – 6
(d) y = –5/3x + 6
-------------
These are all in the form: y = mx + b (m = the slope)
Parallel lines have the same slope, so it's (b) and (c) with a slope of 3/5.

Answer by jim_thompson5910(35256)

Let y=ln(3). So this means that e^y=3


So

e^(ln(3))=e^y=3

So the answer is 3. I don't see "3" as an answer. So make sure that you wrote the problem correctly.